If you have 88gms of C02 AND 36 GRAMS OF H20 REACT, HOW MUCH GLUCOSE AND HOW MUCH OXYGEN ARE PRODUCED ?
1
Expert's answer
2018-10-16T05:35:09-0400

Answer:

264 108 180. 192

6CO2+6H2O= C6H12O6+6O2

88 g. 36 g. X g. Y g.

There is CO2 and H2O equivalent.

m=180 x 88 / 264= 60 g. (C6H12O6)

m= 192x88/264=64 g. (O2)